# Role-Selection Game in Block Production under Proposer-Builder Separation

### v1

@@ Abstract
To address the risks of validator centralization, the Ethereum community introduced Proposer-Builder Separation (PBS), which divides the roles of block building and block proposing to foster a more equitable environment for blockchain participants. PBS creates a two-sided market, wherein searchers provide valuable bundles with bids to builders with the demand for their inclusion in a block, and builders vie for order flows from searchers to secure victory in the block-building auction. For a participant with profit opportunities, strategically selecting their role (either as a searcher or builder) results in varying payoffs. The existence of conflicts or complementarities among bundles leads to an intricate landscape within the PBS ecosystem. In this work, we propose a novel co-evolutionary framework to analyze the behavior of participants in the aforementioned two-sided market, where agents either submit valuable bundles to the private Remote Procedure Call (RPC) of builders or participate in the block-building auction. Agents in our model optimize their strategies through genetic algorithms, grounded in the principles of reinforcement learning. Leveraging agent-based simulations enables us to observe the strategy evolution results of autonomous agents and understand how each profit-seeking actor can benefit from the block-building process under different market conditions. We observe that searchers and builders can develop distinct bidding and rebate strategies under varying conditions (conflict probabilities between bundles), with searchers learning to differentiate their bids based on the rebates offered by different builders. Through empirical game-theoretic analysis, we compute the dynamic equilibrium solution of agents’ strategies under two meta-strategies, which predicts the frequency at which agents employ block building and bundle sharing strategies in the two-sided market. Our analysis reveals that agents achieve a dynamic equilibrium as searchers when the probability of conflict between bundles is low. As this conflict probability rises to a certain critical level, the dynamic equilibrium transitions to favor agents becoming builders.
@@ I Introdution
It is worth noting that MEV is not solely about financial gains and losses; it also has posed a threat to decentralization. The fundamental principle of blockchain technology is that no small group of entities should be able to manipulate the blockchain’s records or impose censorship [4]. In the current Ethereum Proof of Stake (PoS) paradigm, traditional miners in the Proof of Work (PoW) are replaced by validators, and these block producers can allocate MEV rewards as additional stakes to block reward, bolstering their power over the protocol. Thus, the asymmetry in block producers’ capacity to extract MEV constitutes a long-term force driving centralization [5, 6]. As a countermeasure to this trend, the Ethereum community introduced Proposer-Builder Separation (PBS) [7] to ensure a fairer landscape. The logic behind PBS is to split the role of the validators into two parts, separating block building from block proposal. Specialized entities, known as block builders, are responsible for creating the most profitable blocks, while the block proposer (the chosen PoS validator) selects the block with the highest bid through an auction  [^1] Flashbots’ MEV-Boost building auction exemplifies a successful out-of-protocol PBS implementation, with approximately 90% of all new blocks being created via the MEV-Boost block [8]. among builders to propose it on the blockchain. This not only creates opportunities to prevent transaction censorship at the protocol level but also levels the MEV extraction ability between hobbyists and institutional validators.
@@
Despite the introduction of PBS to tackle the issue, it appears ineffective in resolving the problem. Currently, three builders—beaverbuild, rsync, and Titan (BRT)—are responsible for constructing nearly all the blocks [8, 9], which continues to compromise Ethereum’s censorship resistance and decentralization. Moreover, such concentration creates a kind of barrier to entry for new builders, exemplified by a ”chicken-and-egg” problem [8, 10, 9]. Decentralized block building has become the primary improvement target of the current PBS framework, wherein searchers and builders can collaborate to construct the optimal block together by leveraging cryptographic technologies, such as Trusted Execution Environments [11] and Multi-Party Computation [12]. The concept of such a decentralized block building naturally prompts inquiries regarding how each profit-seeking actor can derive benefits from the block-building process. To address this question, it is essential to model the strategies employed by these profit-seeking actors.
@@
In this work, we propose a novel co-evolutionary framework for modeling the block building process. We conceptualize the block building process as a complex adaptive system [13] wherein profit-seeking agents can choose to act as builders, participating in the block building auction under the PBS framework to obtain surplus, or as searchers, sharing a portion of their profit opportunities with other builders to collaboratively construct blocks and avoid involvement in the block building auction. Each agent continuously optimizes its strategy through a genetic algorithm based on the principles of reinforcement learning. Through agent-based simulation, we investigated the co-evolutionary dynamics of agents’ strategies, establishing the feasibility of our co-evolutionary framework.
@@
    We consider a two-sided market in our model, inherently more challenging to manage than a one-sided market, involving bundle interactions as evidenced in the block building process utilizing a greedy algorithm. In this model, agents’ strategies are not fixed arbitrarily; they emerge from a learning process based on a co-evolution supported by genetic algorithms. (see Sect. II-B for details).
@@
    Our agent-based simulations enable us to clearly observe the co-evolutionary outcomes of the agents (see Section III-A for details). In addition, we analyze the impact of bundle interactions on the evolution of agents’ strategies and the distribution of benefits among participants. Our findings reveal a non-monotonic effect of the conflict probability between bundles on the co-evolutionary outcomes of the system (see Sect. III-B for details).
@@
    We consider the role selection of agents in the block building process as a meta-game, which encompasses meta-strategies of block building and bundle sharing. Based on this, we conduct an empirical game-theoretic analysis [14] and calculate the dynamic equilibrium solutions of agents’ strategies using the $\alpha$-Rank [15] method, enabling us to predict the landscape of the block building process across different market conditions (see Sect. III-C for details).
II Model
In this section, we formalize the agent-based model. Unlike the legacy blockchain system, where validators function as both block proposers and validators, in the PBS ecosystem depicted in Fig. 1, more specialized Private Order Flow (POF) providers - searchers, send valuable bundles containing their own transactions and potentially other transactions from the Ethereum mempool, along with their bid for inclusion, to one or multiple reputable block builders’ private Remote Procedure Call (RPC) endpoints instead of the public mempool, in order to conceal their own arbitrage opportunities. For block builders, securing victory in the block building auction hinges on obtaining differentiated POF, which is only achievable if they possess a significant market share.
Figure 1: Sketch for searchers sending bundles to a builder via private channel.
@@ II Model
Let $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{n}\}$, where $n>2$, represent a collection of agents participating in block building process. We represent the agents’ index set by $\mathcal{N}=\{1,2,\cdots,n\}$. Each agent $a_{i}$ is associated with a unique arbitrage opportunity defined as a bundle $\mathbf{b}_{i}$. The bundle must be included in a block $\mathcal{B}_{j}$ and land on the chain to effectively extract value. A block is an ordered tuple $(\mathbf{b}_{(1)},\mathbf{b}_{(2)},\cdots,\mathbf{b}_{(m)})$ composed of bundles, where the index indicates the execution order of the bundle within the block. Within the bundle, transactions run as smart contracts to seize arbitrage opportunities in the market, while potentially being impacted by other bundles in the same block. We signify the effective value of the bundle $\mathbf{b}_{i}$ in block $\mathcal{B}_{j}$ as $v(\mathbf{b}_{i},\mathcal{B}_{j})$.
@@
II-A Agent Policies
In our model, we transform the roles of searchers and builders into two strategies employed by all agents: bundle sharing and block building.
As reported by [10], searchers can employ a ”shotgun” tactic, meaning they submit to more than four builders or even all known builders. Such searcher submission preferences are presumed by [6] as contributing to the utility of a prototypical order flow provider. Our model relaxes the relevant assumptions by considering agents as risk-neutral profit seekers, whose behavior results from strategic co-evolution in a two-sided market.
Bundle sharing strategy involves a bidding vector $\mathbf{p}_{i}=(p_{i1},p_{i2},\cdots,p_{in})$ of agent $i$, where $p_{ij}:=p_{ij}(v(\mathbf{b}_{i},\mathcal{B}_{j}))$ represents the bid agent $i$ is willing to pay agent $j$ for inclusion of bundle $\mathbf{b}_{i}$ in the block $\mathcal{B}_{j}$. In addition, [16] found that the most commonly used strategy involves bribing a certain proportion of profit, which means $p_{ij}(v(\mathbf{b}_{i},\mathcal{B}_{j}))=\beta_{ij}\cdot v(\mathbf{b}_{i},\mathcal{B}_{j}),\beta_{ij}\in(0,1)$. Even the top-earning searchers utilize this method due to its simplicity. It allows the bid to adjust linearly with profit, ensuring that the bid does not exceed the profit during the block building phase. Bundle building as a strategy represents a rebate ratio $\alpha_{j}\in[0,1)$ of agent $j$, where means the proportion of builder surplus refunded to searcher.
@@ II-A Agent Policies
$\pi_{j}(\alpha_{j})=\begin{cases}(1-\alpha_{j})\left(\sum_{i\in\mathcal{N}}p_{ij}-P_{j}\right),&\mathcal{B}_{j}=\mathcal{B}^{*},\\ 0,&\mathcal{B}_{j}\neq\mathcal{B}^{*}\end{cases}$
@@
In a one-sided market consisting of one searcher and two builders, a rational searcher will ignore the builders’ rebate ratio signals and consistently opt to submit bids of zero, prompting the builders to subsequently select a refund of zero for the surplus. This is based on the derivation of a Bayesian Nash equilibrium in an extensive game, which we provide details in Appendix A. However, in a two-sided market that incorporates randomly interacting bundles, the situation becomes more complex and challenging to analyze directly, which will be the main focus of our subsequent discussion.
II-B Modeling of Co-evolution

### v2

@@ Abstract
To address the risks of validator centralization, Proposer-Builder Separation (PBS) was introduced in Ethereum to divide the roles of block building and block proposing, fostering a more equitable and decentralized block production environment. PBS creates a two-sided market in which searchers submit valuable bundles to builders for inclusion in blocks, while builders compete in auctions for block proposals. In this paper, we formulate and analyze a role-selection game that models how profit-seeking participants in PBS strategically choose between acting as searchers or builders, using a co-evolutionary framework to capture the complex interactions and payoff dynamics in this market. Through agent-based simulations, we demonstrate that agents’ optimal role—acting as searcher or builder—responds dynamically to the probability of conflict between bundles. Our empirical game-theoretic analysis quantifies the equilibrium frequencies of role selection under different market conditions, revealing that low conflict probabilities lead to equilibria dominated by searchers, while higher probabilities shift equilibrium toward builders. Additionally, bundle conflicts have non-monotonic effects on agent payoffs and strategy evolution. Our results advance the understanding of decentralized block building and provide guidance for designing fairer and more robust block production mechanisms in blockchain systems.
@@ I Introdution
It is worth noting that MEV is not solely about financial gains and losses; it also has posed a threat to decentralization. The fundamental principle of blockchain technology is that no small group of entities should be able to manipulate the blockchain’s records or impose censorship [4]. In the current Ethereum Proof of Stake (PoS) paradigm, traditional miners in the Proof of Work (PoW) are replaced by validators, and these block producers can allocate MEV rewards as additional stakes to block reward, bolstering their power over the protocol. However, the asymmetry in block producers’ capacity to extract MEV constitutes a long-term force driving centralization [5, 6]. As a countermeasure to this trend, the Ethereum community introduced Proposer-Builder Separation (PBS) [7] to ensure a fairer landscape. The logic behind PBS is to split the role of the validators into two parts, separating block building from block proposal. Specialized entities, known as block builders, are responsible for creating the most profitable blocks, while the block proposer (the chosen PoS validator) selects the block with the highest bid through an auction [^1] Flashbots’ MEV-Boost building auction exemplifies a successful out-of-protocol PBS implementation, with approximately 90% of all new blocks being created via the MEV-Boost block [8]. among builders to propose it on the blockchain. This not only creates opportunities to prevent transaction censorship at the protocol level but also levels the MEV extraction ability between hobbyists and institutional validators.
@@
Despite the introduction of PBS to tackle the issue, it appears ineffective in resolving the problem. Currently, three builders—beaverbuild, rsync, and Titan (BRT)—are responsible for constructing nearly all the blocks [8, 9], which continues to compromise Ethereum’s censorship resistance and decentralization. Moreover, such concentration creates a kind of barrier to entry for new builders, exemplified by a “chicken-and-egg” problem [8, 10, 9]. Decentralized block building has become the primary improvement target of the current PBS framework, wherein searchers and builders can collaborate to construct the optimal block together by leveraging cryptographic technologies, such as Trusted Execution Environments [11] and Multi-Party Computation [12]. The concept of such a decentralized block building naturally prompts inquiries regarding how each profit-seeking actor can derive benefits from the block-building process. To address this question, it is essential to model the strategies employed by these profit-seeking actors.
@@
In this work, we study a strategic interaction wherein agents endowed with profit opportunities can choose between two roles: acting as searchers, who share a portion of their profit opportunities with builders to collaboratively construct blocks and thereby avoid involvement in the block building auction, or as builders, who participate in the auction under the PBS framework in pursuit of surplus. We model this block building game as a complex adaptive system [13], where agents continuously optimize their strategies using a genetic algorithm inspired by reinforcement learning principles. Through agent-based simulation, we investigate the co-evolutionary dynamics of agents’ strategies, demonstrating the feasibility and rich behavioral implications of our proposed framework. [^2] The code supporting the findings of this paper is available at https://github.com/EthanLIYanzhen/Role-Selection-Game-in-Block-Production-under-Proposer-Builder-Separation.git
@@
    We consider a two-sided market in our model, inherently more challenging to manage than a one-sided market, involving bundle interactions as evidenced in the block building process utilizing a greedy algorithm. In this model, agents’ strategies are not fixed arbitrarily; they emerge from a learning process based on a co-evolution supported by genetic algorithms.
@@
    Our agent-based simulations enable us to clearly observe the co-evolutionary outcomes of the agents. In addition, we analyze the impact of bundle interactions on the evolution of agents’ strategies and the distribution of benefits among participants. Our findings reveal a non-monotonic effect of the conflict probability between bundles on the co-evolutionary outcomes of the system.
@@
    We consider the role selection of agents in the block building process as a meta-game, which encompasses meta-strategies of block building and bundle sharing. Based on this, we conduct an empirical game-theoretic analysis (EGTA) [14] and calculate the dynamic equilibrium solutions of agents’ strategies using the $\alpha$-Rank [15] method, enabling us to predict the landscape of the block building process across different market conditions.
II Related Works
II-A Transaction Fee Mechanisms
Certain researchers employ mechanism design theory to evaluate the feasibility of establishing a dominant-strategy incentive-compatible transaction fee mechanism [16, 17, 18, 19, 20]. However, the impossibility results derived from these theoretical studies indicate that the desired attributes of an ideal transaction fee mechanism cannot be achieved simultaneously. For instance, Bahrani et al. [19] illustrate that no fee mechanism can attain incentive compatibility if block producers strive to actively extract MEV, as commonly observed in practice. The analysis in [18], focused on a transaction fee mechanism (TFM) problem for a model featuring searchers and block producers (builder-proposer integration in PBS), considers a more detailed block production process, particularly a decentralized version of PBS. Unlike the literature grounded in mechanism design, this paper primarily focuses on the strategies of participants within the MEV supply chain under the more practical framework of non-revealing mechanisms, specifically the first-price auction.
II-B Empirical Studies on PBS
In recent literature, several studies have investigated the PBS landscape, underscoring the mounting tendency towards centralization of block building within the framework of PBS. Heimbach et al. [21] provide a comprehensive empirical analysis of Ethereum’s PBS adoption, revealing persistent centralization among builders and relays, while also highlighting that PBS, despite intentions, has not effectively mitigated censorship and often relies on the trustworthiness of relays. Gupta et al. [22] empirically demonstrate that builders proficient in capturing CEX/DEX arbitrage (e.g., HFT firms) can construct blocks with significant top-of-block value. Yang et al. [9] propose two metrics to quantify the competitiveness and efficiency of the MEV-boost auction, highlighting issues related to current PBS, including entry barriers for builders and trust crises faced by searchers. Öz et al. [8] reveal that builders’ profitability is correlated with access to exclusive order flow, which is subsequently linked to their market share, thereby naturally underscoring a “chicken-and-egg” dilemma.
II-C Game-Theoretic Approaches to PBS
Some scholars have performed theoretical analyses of the PBS framework using game-theoretic models. Bahrani et al. [5] assess the effectiveness of PBS, quantifying how the competitiveness of the builder ecosystem influences the ability of PBS to mitigate reward heterogeneity among proposers. In addition to their empirical work discussed earlier, Gupta et al. [22] also develop a game-theoretic model illustrating how builders with superior top-of-block capabilities can reinforce their dominance in order flow and PBS auctions through a positive feedback mechanism. Capponi et al. [6] propose a three-stage game-theoretical model involving heterogeneous builders and a single order flow provider, further elucidating the mechanisms underlying builder dominance. Additionally, Mamageishvili et al. [23] examine how competition among searchers influences the allocation of value between a single validator and multiple searchers. Wu et al. [24, 25] investigate strategic behavior among builders in MEV-Boost auctions under the PBS framework, utilizing agent-based simulations and EGTA. They analyze various bidding strategies and latency conditions, finding that a decentralized builder market often incentivizes collusion, resulting in suboptimal auction outcomes. Braga et al. [26] extend the analysis of MEV and PBS by proposing a dynamic protocol-level MEV sharing mechanism modeled as an adjustable extraction rate. They analyze its evolution and stability through dynamical systems theory, demonstrating that suitable adjustment intensities yield a stable equilibrium between users and block producers. Even under chaotic dynamics, the system maintains bounded deviations from equilibrium and ensures long-term market liveness, thus robustly guaranteeing MEV sharing.
III Model
In the PBS ecosystem illustrated in Fig. 1, specialized Private Order Flow (POF) providers—known as searchers—send valuable bundles containing their own transactions and bids for inclusion directly to block builders via private channels, bypassing the public mempool to protect their strategies and arbitrage opportunities. Block builders aggregate these differentiated order flows and participate in the block building auction.
Figure 1: Sketch of the PBS ecosystem
@@ III Model
Let $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{n}\}$, where $n>2$, represent a collection of agents participating in block building process. We represent the agents’ index set by $\mathcal{N}=\{1,2,\cdots,n\}$. Each agent $a_{i}$ is associated with a unique arbitrage opportunity defined as a bundle $\mathbf{b}_{i}$. The bundle must be included in a block $\mathcal{B}_{j}$ and land on the chain to effectively extract value. A block is an ordered tuple $(\mathbf{b}_{(1)},\mathbf{b}_{(2)},\cdots,\mathbf{b}_{(m)})$ composed of bundles, where the index indicates the execution order of the bundle within the block. Within a bundle, transactions are executed sequentially to capture profit opportunities in the market, but their outcomes may be influenced by the presence and ordering of other bundles within the same block. We signify the effective value of the bundle $\mathbf{b}_{i}$ in block $\mathcal{B}_{j}$ as $v(\mathbf{b}_{i},\mathcal{B}_{j})$.
@@
III-A Agent Policies
In our model, we transform the roles of searchers and builders into two strategies employed by all agents: bundle sharing and block building. [^3] As reported by [10], searchers can employ a “shotgun” tactic, meaning they submit to more than four builders or even all known builders. Such searcher submission preferences are presumed by [6] as contributing to the utility of a prototypical order flow provider. Our model relaxes the relevant assumptions by considering agents as risk-neutral profit seekers, whose behavior results from strategic co-evolution in a two-sided market. Bundle sharing strategy involves a bidding vector $\mathbf{p}_{i}=(p_{i1},p_{i2},\cdots,p_{in})$ of agent $i$, where $p_{ij}:=p_{ij}(v(\mathbf{b}_{i},\mathcal{B}_{j}))$ represents the bid agent $i$ is willing to pay agent $j$ for inclusion of bundle $\mathbf{b}_{i}$ in the block $\mathcal{B}_{j}$. In addition, [27] found that the most commonly used strategy involves bribing a certain proportion of profit, which means $p_{ij}(v(\mathbf{b}_{i},\mathcal{B}_{j}))=\beta_{ij}\cdot v(\mathbf{b}_{i},\mathcal{B}_{j}),\beta_{ij}\in(0,1)$. Even the top-earning searchers utilize this method due to its simplicity. It allows the bid to adjust linearly with profit, ensuring that the bid does not exceed the profit during the block building phase. Bundle building as a strategy represents a rebate ratio $\alpha_{j}\in[0,1)$ of agent $j$, where means the proportion of builder surplus refunded to searcher.
@@ III-A Agent Policies
$\pi_{j}(\alpha_{j})=\begin{cases}(1-\alpha_{j})\left(\sum_{i\in\mathcal{N}}p_{ij}-P_{j}\right),&\mathcal{B}_{j}=\mathcal{B}^{*},\\ 0,&\mathcal{B}_{j}\neq\mathcal{B}^{*}.\end{cases}$
In a one-sided market with one searcher and two builders, a rational searcher will disregard the builders’ rebate ratio signals and always choose to submit a bid of zero, leading the builders to in turn set the surplus refund to zero. This follows from the analysis of the Bayesian Nash equilibrium in the corresponding extensive game, the details of which are presented below. Specifically, consider a setting with $|\mathcal{A}|=3$, where agents $a_{1}$ and $a_{2}$ serve as builders and $a_{3}$ acts as the searcher. Due to severe conflict between the builders, with $\varphi_{1,2}=\varphi_{2,1}=-1$, they cannot share bundles with each other. We further assume that $a_{3}$ always adopts the bundle-sharing strategy. Each agent’s bundle value $v_{i}$ is independently drawn from a known distribution $F_{i}$. Agents do not observe each other’s private valuations, but the value distributions are common knowledge.
The payoff functions $\pi_{1}$, $\pi_{2}$, and $\pi_{3}$ are defined according to the values of the corresponding variables and can be summarized as follows.
For agent $a_{1}$, the payoff function is given by $\pi_{1}(v_{2},v_{3})=(1-\alpha_{13})[v_{1}-v_{2}+(\beta_{31}-\beta_{32})v_{3}]$ if $v_{1}+\beta_{31}v_{3}\geq v_{2}+\beta_{32}v_{3}$; otherwise, $\pi_{1}(v_{2},v_{3})=0$.
Similarly, for agent $a_{2}$, when $v_{2}+\beta_{32}v_{3}\geq v_{1}+\beta_{31}v_{3}$, we have $\pi_{2}(v_{1},v_{3})=(1-\alpha_{23})[v_{2}-v_{1}+(\beta_{32}-\beta_{31})v_{3}]$, and $\pi_{2}(v_{1},v_{3})=0$ otherwise.
For agent $a_{3}$, the payoff is $\pi_{3}(v_{1},v_{2})=(1-\beta_{31})v_{3}+\alpha_{13}[v_{1}-v_{2}+(\beta_{31}-\beta_{32})v_{3}]$ if $v_{1}+\beta_{31}v_{3}\geq v_{2}+\beta_{32}v_{3}$, and $\pi_{3}(v_{1},v_{2})=(1-\beta_{32})v_{3}+\alpha_{23}[v_{2}-v_{1}+(\beta_{32}-\beta_{31})v_{3}]$ otherwise.
Denote $\mathrm{\Delta}v=v_{1}-v_{2}$ and $\mathrm{\Delta}\beta=\beta_{31}-\beta_{32}$. Then, the payoff function $\pi_{3}(\mathrm{\Delta}v)$ can be written as follows: when $\mathrm{\Delta}v\geq-\mathrm{\Delta}\beta\,v_{3}$, we have $\pi_{3}(\mathrm{\Delta}v)=v_{3}-(\beta_{32}+\mathrm{\Delta}\beta)\,v_{3}+\alpha_{13}(\mathrm{\Delta}v+\mathrm{\Delta}\beta\,v_{3})$; otherwise, i.e., when $\mathrm{\Delta}v<-\mathrm{\Delta}\beta\,v_{3}$, the payoff becomes $\pi_{3}(\mathrm{\Delta}v)=v_{3}-\beta_{32}v_{3}+\alpha_{23}\big{(}-\mathrm{\Delta}v-\mathrm{\Delta}\beta\,v_{3}\big{)}$.
$f_{\mathrm{\Delta}v}(x)=\begin{cases}\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{-\lambda_{1}x},&x\geq 0,\\ \frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{\lambda_{2}x},&x<0.\end{cases}$
Accordingly, the expected payoff $\mathbb{E}[\pi_{3}]$ can be expressed as
$\mathbb{E}[\pi_{3}]$
$=$ $\int_{-\mathrm{\Delta}\beta v_{3}}^{+\infty}[v_{3}-(\beta_{32}+\mathrm{\Delta}\beta)v_{3}+\alpha_{13}(x+\mathrm{\Delta}\beta v_{3})]f_{\mathrm{\Delta}v}(x)\mathrm{d}x$
$\mathbb{E}[\pi_{3}]$
$=$ $\int_{0}^{+\infty}[v_{3}-\mathrm{\Delta}\beta v_{3}+\alpha_{13}(x+\mathrm{\Delta}\beta v_{3})]\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{-\lambda_{1}x}\mathrm{d}x$
$+\int_{-\infty}^{-\mathrm{\Delta}\beta v_{3}}[v_{3}+\alpha_{23}(-x-\mathrm{\Delta}\beta v_{3})]\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{\lambda_{2}x}\mathrm{d}x.$
Computing the derivative of $\mathbb{E}[\pi_{3}]$ with respect to $\mathrm{\Delta}\beta$, we have
$\frac{\mathrm{d}\mathbb{E}[\pi_{3}]}{\mathrm{d}\mathrm{\Delta\beta}}$
=\,= $\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\Bigg{(}\frac{v_{3}(\alpha_{13}-1)}{\lambda_{1}}+\frac{v_{3}(\alpha_{13}-1)}{\lambda_{2}}\left[1-e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}\right]$
$\hskip 79.6678pt-v_{3}^{2}\,\mathrm{\Delta}\beta\,e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}-\frac{\alpha_{23}v_{3}\,e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}}{\lambda_{2}}\Bigg{)}$
<\,< $0$
@@
This implies that the searcher should adopt a $\mathrm{\Delta}\beta=0$ strategy, meaning uniform bidding to all builders. However, in a two-sided market that incorporates randomly interacting bundles, the situation becomes more complex and challenging to analyze directly, which will be the main focus of our subsequent discussion.
III-B Modeling of Co-evolution